OpenStudy (anonymous):

Simplify: (b3 – 11) – (12b + 4) -11b4 – 15 b3 – 12b – 7 -11b4 – 7 b3 – 12b – 15

OpenStudy (theviper):

=b3-11-12b-4 =b3-12b-15

OpenStudy (theviper):

So last one is correct:)

OpenStudy (anonymous):

Find the perimeter of a rectangle with length 5b – 4 and width 4a2 20a2b + 16a2 20a2b – 16a2 4a2 + 5b – 4 8a2 + 10b – 8

OpenStudy (theviper):

Perimeter of a rect.=2(l+b)

OpenStudy (theviper):

2(5b-4+4a2)

OpenStudy (theviper):

=10b-8+8a2

OpenStudy (theviper):

Clear???

OpenStudy (anonymous):

yea thanks :)

OpenStudy (theviper):

yw:)

OpenStudy (theviper):

In which class do u study?

OpenStudy (theviper):

@ladynisha

OpenStudy (theviper):

?????????

OpenStudy (theviper):

Not want to tell no problem:) It's ur wish.